### v1

# On Wasserstein distances for affine transformations of random vectors

Abstract.
We expound on some known lower bounds of the quadratic Wasserstein distance between random vectors in $\mathbb{R}^{n}$ with an emphasis on affine transformations that have been used in manifold learning of data in Wasserstein space. In particular, we give concrete lower bounds for rotated copies of random vectors in $\mathbb{R}^{2}$ with uncorrelated components by computing the Bures metric between the covariance matrices. We also derive upper bounds for compositions of affine maps which yield a fruitful variety of diffeomorphisms applied to an initial data measure. We apply these bounds to various distributions including those lying on a 1-dimensional manifold in $\mathbb{R}^{2}$ and illustrate the quality of the bounds. Finally, we give a framework for mimicking handwritten digit or alphabet datasets that can be applied in a manifold learning framework.
1. Introduction
Use of optimal transport as a powerful tool in data analysis and machine learning has increased significantly in recent years. The power of the substantial theoretical underpinnings of optimal transport make it particularly appealing, as does its success in the various ways it has been deployed. Of primary importance is the fact that optimal transport gives an optimal (in terms of a prescribed cost function) way of transforming one data measure into another, and indeed the associated Wasserstein distance defines a metric on the space of probability measures with an appropriate number of finite moments. While Euclidean data is abundantly common, oftentimes one can also view data as probability measures in some natural way (e.g., pixel images can be mapped to a uniform grid and assigned a discrete measure based on intensity).
One the many ways optimal transport can be used in machine learning is in manifold learning, which is a technique that assumes data lives on a low-dimensional manifold embedded in a high-dimensional space (in this case $W_{2}(\mathbb{R}^{d})$). Understanding submanifolds of Wasserstein space is the subject of ongoing research [4, 9, 12, 13, 14, 20]. Toward this, several works have considered manifolds generated by translating, dilating, or rotating a single probability measure. Most of these techniques then require computing the Wasserstein distances between the original measure and its transformed copies.
In easy cases like translation and dilation, one can explicitly compute these distance, but for rotations and more sophisticated transformations this is highly non-trivial. There are some known upper and lower bounds for the Wasserstein distance [5, 7] and equality of the lower bound is known to hold for Gaussians [8] and certain structured distributions. Recent manifold learning results [4, 9, 14] show that one can isometrically embed translation and dilation manifolds into Euclidean space, while providing evidence that such embeddings can still sometimes find the structure of a rotational manifold (though this can break down in general [14]).
To better understand rotational submanifolds of Wasserstein space, we expound on the lower bounds in terms of the Bures metric. In particular, for $2$-dimensional rotations of random vectors with uncorrelated components, we give an explicit formula for the lower bound in terms of the means and variances of each component and the rotation angle. We illustrate the sharpness (or not) of the lower bound for rotations in cases when the lower bound is known not to be attained. In many cases, we see that it provides a good estimation of the actual Wasserstein distance.
Additionally, we consider compositions of translations, dilations, and rotations as a way to model more complex transformations of data, and provide an upper bound for some such transformations. Finally, we show how one can generate handwritten alphabet letters on which these transformations can be applied to give rise to a synthetic handwriting dataset that can be used for study of manifold learning algorithms in a controlled setting.
1.1. Notation and definitions
Let $\mu$ be a probability measure on ${\mathbb{R}^{n}}$, and let $y\in\mathbb{R}^{n}$ and $A\in\mathbb{R}^{n\times n}$ be fixed. Let $T:{\mathbb{R}^{n}}\to{\mathbb{R}^{n}}$ be an affine map given by
$T(x)=y+Ax,\quad x\in\mathbb{R}^{n}.$
Then $T$ induces a probability measure as follows:
$\mu_{T}(B):=\mu(T^{-1}(B)),\quad B\in\mathcal{B}(\mathbb{R}^{n}),$
where $\mathcal{B}(\mathbb{R}^{n})$ is the Borel $\sigma$-algebra on $\mathbb{R}^{n}$. In many cases $\mu_{T}$ is called the pushforward of $\mu$ under $T$ and may be denoted $T_{\#}\mu$ (e.g., [18]).
Any probability measure $\mu$ can be viewed as the distribution of a random vector defined on some probability space $(\Omega,\Phi,P)$. Namely, take $(\Omega,\Phi,P)=(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}),\mu)$ and consider
$X:(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}),\mu)\to(\mathbb{R}^{n},\mathcal{B}(\mathbb{R}^{n}),P_{X})$
defined by the identity map $X(x)=x$, where $P_{X}$ is the distribution of $X$ given by
$P_{X}(B)=P\left(\omega\in\Omega\;|\;X(\omega)\in B\right)=\mu\left(x\in\mathbb{R}^{n}:X(x)\in B\right)=\mu(X^{-1}(B))=\mu(B).$
In this case, $\mu$ is often called the law of the random vector $X$.
In other words, for any probability measure $\mu$ on $\mathbb{R}^{n}$, there exists a random vector $X\in\mathbb{R}^{n}$ whose probability distribution $P_{X}$ coincides with $\mu$. Consequently, one can describe measure transformation $\mu\mapsto\mu_{T}$ in terms of the distribution of the corresponding random vectors $X\mapsto T(X)$ by
$P_{X}\mapsto P_{T(X)},\quad\textnormal{where}\quad P_{T(X)}(B)=P_{X}(T^{-1}(B)).$
Two salient properties of random vectors are their mean and covariance matrices defined by
$m_{X}=\mathbb{E}[X],\quad\textnormal{and}\quad\Sigma_{X}=\mathbb{E}\left[(X-m_{X})(X-m_{X})^{\top}\right],$
respectively. We say that a random vector $X\in\mathbb{R}^{n}$ with components $X_{1},\dots,X_{n}$ has uncorrelated components if $\mathbb{E}[X_{i}X_{j}]=0$ for all $i\neq j$. The variance of a component is $\operatorname{Var}[X_{i}]=\mathbb{E}[X_{i}^{2}]-\mathbb{E}[X_{i}]^{2}$.
In what follows, we express the quadratic Wasserstein distance between two measures $\mu$ and $\nu$ in terms of their corresponding random vectors $X$ and $Y$ via
$W_{2}(\mu,\nu):=W_{2}(X,Y):=\min\left\{(\mathbb{E}|X-Y|^{2})^{\frac{1}{2}}:X\sim\mu,Y\sim\nu\right\},$ (1)
where $X\sim\mu$ means that the law of $X$ is $\mu$, and similarly for $Y\sim\nu$. Some basic properties are that $W_{2}$ defines a metric on the subset of $\mathcal{P}(\mathbb{R}^{n})$ of measures with finite second moment, and is intimately related to optimal transport theory, which has found application in a vast array of fields including PDEs, signal processing, and machine learning, to name a very few. For seminal works on optimal transport theory, see [18, 19], and for more applied and computational treatments [16, 17].
Proposition 1.1.
Given $\mu,\nu\in\mathcal{P}(\mathbb{R}^{n})$ with corresponding random vectors $X,Y\in\mathbb{R}^{n}$, we have
$W_{2}(\mu,\nu)^{2}=|m_{X}-m_{Y}|^{2}+\min_{X\sim\mu,Y\sim\nu}\mathbb{E}|(X-m_{X})-(Y-m_{Y})|^{2}.$ (2)
Proof.
Note that
$\mathbb{E}|X-Y|^{2}$ $=\mathbb{E}|(X-m_{X})-(Y-m_{Y})+(m_{X}-m_{Y})|^{2}$
$=\mathbb{E}|m_{X}-m_{Y}|^{2}+\mathbb{E}|(X-m_{X})-(Y-m_{Y})|^{2}+2\mathbb{E}\left[\left((X-m_{X})-(Y-m_{Y})\right)(m_{X}-m_{Y})\right]$
$=|m_{X}-m_{Y}|^{2}+\mathbb{E}|(X-m_{X})-(Y-m_{Y})|^{2}+2(m_{X}-m_{Y})\left(\mathbb{E}[X-m_{X}]-\mathbb{E}[Y-m_{y}]\right)$
$=|m_{X}-m_{Y}|^{2}+\mathbb{E}|(X-m_{X})-(Y-m_{Y})|^{2},$
where the final equality comes from the fact that $\mathbb{E}[X-m_{X}]=\mathbb{E}[Y-m_{Y}]=0$. The result follows from putting the right hand side above into (1). ∎
The utility of the above proposition is that computing the quadratic Wasserstein distance between two random vectors requires only computing the minimizer of the second term in (2).
1.2. Lower bounds for the Wasserstein distance
Proposition 1.1 implies that to better understand the Wasserstein distance, one needs to understand $\mathbb{E}|X-Y|^{2}$ for zero-mean $X$ and $Y$. To that end, we invoke the following result of Dowson and Landau; note that they call $W_{2}$ the Fréchet distance, which appears to be more widely used in probability theory.
Theorem 1.2 ([5]).
Let $X,Y$ be random vectors in $\mathbb{R}^{n}$ having zero mean and covariance matrices $\Sigma_{X}$ and $\Sigma_{Y}$, respectively. Then
$\mathbb{E}|X-Y|^{2}\geq\textnormal{tr}[\Sigma_{X}+\Sigma_{Y}-2(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}]$
where the square roots are the positive roots. Moreover, the bound is attained when $X-Y$ has covariance matrix
$\Sigma_{X-Y}=\Sigma_{X}+\Sigma_{Y}-[(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}+(\Sigma_{Y}\Sigma_{X})^{\frac{1}{2}}].$
If $\Sigma_{X}$ is non-singular, then this occurs when $X$ and $Y$ are related by
$Y=\Sigma_{X}^{-1}(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}X.$
Dowson and Landau proved an upper bound for $\mathbb{E}|X-Y|^{2}$ in which the -- is replaced with $+$ inside the trace formula. This is typically less useful as it is often far from a sharp bound as will be seen in the examples later in the paper. Most of our work is focused on illustrating the quality of the lower bound, hence we focus on in the theorem above. Note also that $B(X,Y):=\textnormal{tr}[\Sigma_{X}+\Sigma_{Y}-2(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}]$ defines a metric (called the Bures metric [2, 6]) on the space of covariance matrices, which is identical to the space of symmetric positive semidefinite matrices.
Now, combining Proposition 1.1 and Theorem 1.2 yields the following.
Corollary 1.3.
Let $\mu,\nu\in\mathcal{P}(\mathbb{R}^{n})$ with corresponding random vectors $X,Y\in\mathbb{R}^{n}$ having means $m_{X},m_{Y}$ and covariance matrices $\Sigma_{X},\Sigma_{Y}$. Then,
$W_{2}(\mu,\nu)^{2}=W_{2}(X,Y)^{2}\geq|m_{X}-m_{Y}|^{2}+\textnormal{tr}[\Sigma_{X-{m_{X}}}+\Sigma_{Y-{m_{Y}}}-2(\Sigma_{X-{m_{X}}}\Sigma_{Y-{m_{Y}}})^{\frac{1}{2}}].$ (3)
While Corollary 1.3 follows from the result of Dowson and Landau, Gelbrich [7] proves a similar result and gives more conditions under which the lower bound is attained (see also the book of Peyré and Cuturi [16, Remarks 2.31 and 2.32]). In particular, Gelbrich proves that the lower bound of Corollary 1.3 is attained if
$\bar{\nu}\circ P_{\mu}^{-1}=\bar{\mu}\circ(\Sigma_{X}^{-\frac{1}{2}}(\Sigma_{X}^{-\frac{1}{2}}(\Sigma_{X}^{\frac{1}{2}}\Sigma_{Y}\Sigma_{X}^{\frac{1}{2}})^{\frac{1}{2}}\Sigma_{X}^{-\frac{1}{2}})^{-1},$ (4)
where $\bar{\mu}$ is the centered version of $\mu$ and $P_{\mu}$ is the orthogonal projection of $\mathbb{R}^{n}$ onto the image (column space) of $\Sigma_{X}$. This condition can be difficult to check, and thus a further result of Gelbrich in the same paper proves that this condition is satisfied by so-called elliptically contoured distributions. An absolutely continuous measure $\mu$ with density $f$ is in this family if, for any $x$, one has
$f(x)=\frac{1}{\sqrt{\det(A)}}h\left(\left\langle x-m_{X},A^{-1}(x-m_{X})\right\rangle\right),$
[…]
Prior to Gelbrich’s work, it was known that equality holds for the lower bound of Corollary 1.3 when $\mu$ and $\nu$ are both Gaussian measures [5, 8]. The arguments of these papers essentially use an optimization result of Olkin and Pukelsheim [15].
It should be noted that in most references other than [5], the trace norm on covariance matrices is expressed as $\textnormal{tr}[\Sigma_{X}+\Sigma_{Y}-2(\Sigma_{X}^{\frac{1}{2}}\Sigma_{Y}\Sigma_{X}^{\frac{1}{2}})^{\frac{1}{2}}]$. We note that these are in fact the same.
Proposition 1.4.
Let $A,B\in\mathbb{R}^{n\times n}$ be symmetric positive semi-definite matrices. Then
$\textnormal{tr}[(AB)^{\frac{1}{2}}]=\textnormal{tr}[(A^{\frac{1}{2}}BA^{\frac{1}{2}})^{\frac{1}{2}}].$
Proof.
Note that $A^{\frac{1}{2}}BA^{\frac{1}{2}}$ is symmetric, positive semidefinite and has the same eigenvalues as $AB=A^{\frac{1}{2}}A^{\frac{1}{2}}B$. Therefore, the square roots of these matrices have the same eigenvalues. The conclusion follows from the fact that the trace of a square matrix is the sum of its eigenvalues. ∎
2. Examples of Wasserstein distances between transformations of a fixed random vector
We now undertake the task of computing the Wasserstein distance between a random vector and various simple transformations of it explicitly. The examples here are motivated by simple articulations that might arise in an image analysis setting, and which one may want to mod out by (as in group invariant machine learning) or understand an efficient parametrization of (as in manifold learning) depending on the application. The first two examples, in which we explicitly compute the Wasserstein distance between a random vector and its translates or its (anisotropic) dilations are well-known, but our purpose here is to illustrate how both terms of Proposition 1.1 contribute to the Wasserstein distance in these cases.
For translations and dilations of a fixed random vector, equality is achieved in (3) as will be seen below. However, computing the Wasserstein distance between $X$ and a rotated copy of $X$ is unexpectedly tricky. Indeed, Brenier’s celebrated theorem [1] implies that the rotation matrix cannot be the optimal transport map from $X$ to its rotated copy because it is not the gradient of a convex function. In our third example, we compute the lower bound of Corollary 1.3 explicitly for rotations in $\mathbb{R}^{2}$ for zero-mean random vectors with uncorrelated components, and illustrate in subsequent sections how close this bound is in some instances to the actual Wasserstein distance.
2.1. Translations
Given $\alpha\in\mathbb{R}^{n}$, define the translation operator via
$T_{\alpha}(x)=x+\alpha.$
Proposition 2.1.
Let $\mu\in\mathcal{P}(\mathbb{R}^{n})$ with corresponding random vector $X\in\mathbb{R}^{n}$, and let $\Theta\subset\mathbb{R}^{n}$ be a set of translation vectors with corresponding translation operators $\{T_{\theta}:\theta\in\Theta\}$. For any $\theta,\theta^{\prime}\in\Theta$,
$W_{2}(T_{\theta}X,T_{\theta^{\prime}}X)=|\theta-\theta^{\prime}|.$
Moreover, the lower bound of (3) is attained.
Proof.
Note that
$m_{T_{\theta}X}=\mathbb{E}[X+\theta]=\mathbb{E}[X]+\theta=m_{X}+\theta.$
Consequently, $\mathbb{E}|(X+\theta-m_{X+\theta})-(X+\theta^{\prime}-m_{X+\theta^{\prime}})|^{2}=0$, so $W_{2}(T_{\theta}X,T_{\theta^{\prime}}X)=|m_{X}+\theta-(m_{X}+\theta^{\prime})|=|\theta-\theta^{\prime}|$ by Proposition 1.1.
To see that the lower bound of (3) is attained, note that
$\Sigma_{X+\theta-m_{X+\theta}}=\Sigma_{X+\theta-(m_{X}+\theta)}=\Sigma_{X-m_{X}}$
for any $\theta,$ hence the trace part of the lower bound vanishes and the conclusion follows. ∎
Note that one can also verify directly that translations of random vectors satisfy condition (4), which guarantees that the lower bound of (3) is attained. In the case of translations, we see that the difference in means is the only one that contributes to the Wasserstein distance, whereas the trace term vanishes.
2.2. Dilations
Given $\lambda\in\mathbb{R}^{n}$ with no 00 coordinates, define the scaling operator via
$S_{\lambda}(x)=\textnormal{diag}(\lambda_{1},\dots,\lambda_{n})x.$
Proposition 2.2.
Suppose that $\mu\in\mathcal{P}(\mathbb{R}^{n})$ has associated random vector $X=(X_{1},\dots,X_{n})\in\mathbb{R}^{n}$ with uncorrelated components, and hence diagonal covariance matrix given by $\Sigma_{X-m_{X}}=\textnormal{diag}(\operatorname{Var}[X_{1}],\dots,\operatorname{Var}[X_{n}]).$ Let $\Lambda\subset\mathbb{R}^{n}\setminus$ be a set of scaling parameters such that $\lambda\in\Lambda$ satisfies $\lambda_{i}\neq 0$ for all $i$, and let the associated scaling operators be $\{S_{\lambda}:\lambda\in\Lambda\}$. Then for any $\lambda,\lambda^{\prime}\in\Lambda$,
$W_{2}(S_{\lambda}X,S_{\lambda^{\prime}}X)^{2}=\sum_{i=1}^{n}\mathbb{E}[X_{i}^{2}]|\lambda_{i}-\lambda^{\prime}_{i}|^{2}.$
In particular, the lower bound of (3) is attained.
Proof.
Since $=m_{S_{\lambda}X}=\mathbb{E}[S_{\lambda}X]=(\lambda_{1}\mathbb{E}[X_{1}],\dots,\lambda_{n}\mathbb{E}[X_{n}])$, we have
$m_{S_{\lambda}X}-m_{S_{\lambda^{\prime}}X}=((\lambda_{1}-\lambda^{\prime}_{1})\mathbb{E}[X_{1}],\dots,(\lambda_{n}-\lambda^{\prime}_{n})\mathbb{E}[X_{n}]).$
Thus,
$|m_{S_{\lambda}X}-m_{S_{\lambda^{\prime}}X}|^{2}=\sum_{i=1}^{n}\mathbb{E}[X_{i}]^{2}|\lambda_{i}-\lambda^{\prime}_{i}|^{2}.$
Now we compute the minimum over $X$ of $\mathbb{E}[|(S_{\lambda}X-m_{S_{\lambda}X})-(S_{\lambda^{\prime}}X-m_{S_{\lambda^{\prime}}X})|^{2}]$ as follows.
$\mathbb{E}[|(S_{\lambda}X-m_{S_{\lambda}X})-(S_{\lambda^{\prime}}X-m_{S_{\lambda^{\prime}}X})|^{2}]$ $=\mathbb{E}[|(\lambda_{i}-\lambda_{i}^{\prime})X_{i})_{i}-((\lambda_{i}-\lambda_{i}^{\prime})\mathbb{E}[X_{i}])_{i}|^{2}]$
$=\mathbb{E}[|((\lambda_{i}-\lambda_{i}^{\prime})(X_{i}-\mathbb{E}[X_{i}]))_{i}|^{2}]$
$=\sum_{i=1}^{n}|\lambda_{i}-\lambda_{i}^{\prime}|^{2}\mathbb{E}[|X_{i}-\mathbb{E}[X_{i}]|^{2}]$
$=\sum_{i=1}^{n}|\lambda_{i}-\lambda_{i}^{\prime}|^{2}(\mathbb{E}[X_{i}^{2}]-\mathbb{E}[X_{i}]^{2}).$
Consequently,
$W_{2}(S_{\lambda}X,S_{\lambda^{\prime}}X)^{2}=\sum_{i=1}^{n}(\mathbb{E}[X_{i}]^{2}+\mathbb{E}[X_{i}^{2}]-\mathbb{E}[X_{i}]^{2})|\lambda_{i}-\lambda_{i}^{\prime}|^{2}=\sum_{i=1}^{n}\mathbb{E}[X_{i}^{2}]|\lambda_{i}-\lambda_{i}^{\prime}|^{2}$
as claimed.
Finally, to see that this is attained via the lower bound of (3), note that one has $\Sigma_{S_{\lambda}X}=\textnormal{diag}(\lambda_{1}^{2}\operatorname{Var}[X_{1}],\dots,\lambda_{n}^{2}\operatorname{Var}[X_{n}]).$ Hence,
$\textnormal{tr}\left(\Sigma_{S_{\lambda}X}+\Sigma_{S_{\lambda^{\prime}}X}-2\left(\Sigma_{S_{\lambda}X}\Sigma_{S_{\lambda^{\prime}}X}\right)^{\frac{1}{2}}\right)$ $=\sum_{i=1}^{n}\left(\lambda_{i}^{2}+\lambda_{i}^{\prime 2}-2\lambda_{i}\lambda^{\prime}_{i}\right)\operatorname{Var}[X_{i}]$
$=\sum_{i=1}^{n}|\lambda_{i}-\lambda^{\prime}_{i}|^{2}\operatorname{Var}[X_{i}]$
$=\sum_{i=1}^{n}|\lambda_{i}-\lambda_{i}^{\prime}|^{2}(\mathbb{E}[X_{i}^{2}]-\mathbb{E}[X_{i}]^{2}).$
In contrast to the translation case, the trace portion of (3) plays a pivotal role in the Wasserstein distance for dilations, and actually cancels out the term involving the difference in means.
2.3. Rotations
In this section, we consider the Wasserstein distance between probability measures under orthogonal transformations on $\mathbb{R}^{n}$, and give an explicit formula for the lower bound of (3) for rotations in $\mathbb{R}^{2}$. To wit, consider a parametrized family of orthogonal transformations $\{Q_{\theta}:\theta\in\Theta\}\subset O(n)$. For a random vector $X\in\mathbb{R}^{n}$ with mean $m_{X}$ and covariance matrix ${\Sigma_{X}}$ we have
${m_{{Q_{\theta}}X}}=\mathbb{E}[Q_{\theta}X]=Q_{\theta}\mathbb{E}[X]=Q_{\theta}m_{X},$
whence, by cyclic properties of trace,
$\textnormal{tr}(\Sigma_{{Q_{\theta}}X})=\textnormal{tr}({Q_{\theta}}{\Sigma_{X}}Q_{\theta}^{\top})=\textnormal{tr}(Q_{\theta}^{\top}{Q_{\theta}}{\Sigma_{X}})=\textnormal{tr}({\Sigma_{X}}).$
Consequently, by Corollary 1.3 and the above, we have, for any parameters $\theta,\varphi\in\Theta$,
$W_{2}(Q_{\theta}X,Q_{\varphi}X)^{2}\geq|Q_{\theta}m_{X}-Q_{\varphi}m_{X}|^{2}+\\ \textnormal{tr}[\Sigma_{X-{m_{X}}}+\Sigma_{X-m_{X}}-2(\Sigma_{{Q_{\theta}X}-m_{Q_{\theta}X}}\Sigma_{{Q_{\varphi}X}-m_{Q_{\varphi X}}})^{\frac{1}{2}}].$ (5)
Remark 2.3.
Since $Y$ and $Y+b$ have the same covariance matrix for any random vector $Y$ and constant vector $b$, the trace part in (5) is
$\textnormal{tr}[2{\Sigma_{X}}-2(\Sigma_{{Q_{\theta}X}}\Sigma_{Q_{\varphi}X})^{\frac{1}{2}}].$
The main challenge in computing this trace is computing the positive square root of
$\Sigma_{{Q_{\theta}X}}{\Sigma_{Q_{\varphi}X}}.$
In what follows, we compute a closed form for the lower bound (5) for rotations in ${\mathbb{R}^{2}}$. To compute the trace part, by Remark 2.3, it suffices to consider a base probability measure corresponding to a zero-mean random vector $X\in\mathbb{R}^{2}$ with uncorrelated components. That is, we assume
$\mathbb{E}[X]=\mathbb{E}\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}=\begin{bmatrix}\mathbb{E}[X_{1}]\\ \mathbb{E}[X_{2}]\end{bmatrix}=\begin{bmatrix}0\\ 0\end{bmatrix},$
and
$\operatorname{Cov}(X_{1},X_{2})=\mathbb{E}[(X_{1}-\mathbb{E}[X_{1}])\;(X_{2}-\mathbb{E}[X_{2}])]=\mathbb{E}[X_{1}X_{2}]=0.$
The family of rotations on $\mathbb{R}^{2}$ are
$R_{\theta}=\begin{bmatrix}\cos\theta&-\sin\theta\\ \sin\theta&\cos\theta\end{bmatrix},\qquad\theta\in[0,2\pi),$
which induce transformations of a random vector $X$ given by
$R_{\theta}X=\begin{bmatrix}X_{1}\cos\theta-X_{2}\sin\theta\\ X_{1}\sin\theta+X_{2}\cos\theta\end{bmatrix}.$
To derive an explicit formula for the trace term in (5) between $R_{\theta}X$ and $R_{\varphi}X$, we need several auxiliary facts which we collect below.
Proposition 2.4.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}\in\mathbb{R}^{2}$ be a zero-mean random vector with uncorrelated components. Then the trace of the covariance matrix of the rotation of $X$ is $\theta$-invariant, and we have for any $\theta\in[0,2\pi)$,
$\textnormal{tr}[{\Sigma_{{R_{\theta}}X}}]=\mathbb{E}[X_{1}^{2}]+\mathbb{E}[X_{2}^{2}].$ (6)
Proof.
We have
${\Sigma_{{R_{\theta}}X}}$ $=\mathbb{E}\begin{bmatrix}X_{1}\cos\theta&-X_{2}\sin\theta\\ X_{1}\sin\theta&X_{2}\cos\theta\end{bmatrix}\begin{bmatrix}X_{1}\cos\theta&-X_{2}\sin\theta\\ X_{1}\sin\theta&X_{2}\cos\theta\end{bmatrix}^{\top}$
$=\begin{bmatrix}\mathbb{E}[(X_{1}\cos\theta-X_{2}\sin\theta)^{2}]&\ast\\ \ast&\mathbb{E}[(X_{1}\sin\theta+X_{2}\cos\theta)^{2}]\end{bmatrix},$
which, using the fact that $\mathbb{E}[X_{1}X_{2}]=0$, reduces to
$\begin{bmatrix}\mathbb{E}[X_{1}^{2}]\cos^{2}\theta+\mathbb{E}[X_{2}^{2}]\sin^{2}\theta&\ast\\ \ast&\mathbb{E}[X_{1}^{2}]\sin^{2}\theta+\mathbb{E}[X_{2}^{2}]\cos^{2}\theta\end{bmatrix},$
which has trace $\mathbb{E}[X_{1}^{2}]+\mathbb{E}[X_{2}^{2}]$ as required.
Lemma 2.5 ([11]).
Let $M\in\mathbb{R}^{2\times 2}$ be given by
$M=\begin{bmatrix}A&B\\ C&D\end{bmatrix}.$
Then
$M^{\frac{1}{2}}=\frac{1}{\sqrt{\textnormal{tr}(M)+2\sqrt{\det(M)}}}\begin{bmatrix}A+\sqrt{\det(M)}&B\\ C&D+\sqrt{\det(M)}\end{bmatrix}.$
Proposition 2.6.
Under the assumptions of Proposition 2.4, for $\theta,\varphi\in[0,2\pi)$, we have
$\textnormal{tr}[(\Sigma_{R_{\theta}X}\Sigma_{R_{\varphi}X})^{\frac{1}{2}}]=\sqrt{(\mathbb{E}[X_{1}^{2}]-\mathbb{E}[X_{2}^{2}])^{2}\cos^{2}(\theta-\varphi)+4\mathbb{E}[X_{1}^{2}]\mathbb{E}[X_{2}^{2}]}.$
Proof.
To simplify notation, set $a=\mathbb{E}[X_{1}^{2}],b=\mathbb{E}[X_{2}^{2}].$ Then,
$(\Sigma_{R_{\theta}X}\Sigma_{R_{\varphi}X})_{11}$ $=(a-b)^{2}\cos\theta\cos\varphi\sin\theta\sin\varphi+(a\cos^{2}\theta+b\sin^{2}\theta)(a\cos^{2}\varphi+b\sin^{2}\varphi),$
$(\Sigma_{R_{\theta}X}\Sigma_{R_{\varphi}X})_{12}$ $=(a-b)\cos(\theta-\varphi)(b\cos\varphi\sin\theta+a\cos\theta\sin\varphi),$
$(\Sigma_{R_{\theta}X}\Sigma_{R_{\varphi}X})_{21}$ $=(a-b)\cos(\theta-\varphi)(b\cos\varphi\sin\theta+a\cos\theta\sin\varphi),$
$(\Sigma_{R_{\theta}X}\Sigma_{R_{\varphi}X})_{22}$ $=(a-b)^{2}\cos\theta\cos\varphi\sin\theta\sin\varphi+(b\cos^{2}\theta+a\sin^{2}\theta)(b\cos^{2}\varphi+a\sin^{2}\varphi).$
From here, we utilize Lemma 2.5 and some tedious calculations involving trigonometric reduction formulas to arrive at
$\textnormal{tr}[(\Sigma_{{R_{\theta}}X}\Sigma_{{R_{\varphi}}X})^{\frac{1}{2}}]=\sqrt{{{(a-b)}^{2}}{{\cos}^{2}}(\theta-\varphi)+4ab}=\sqrt{{{(a+b)}^{2}}-{{(a-b)}^{2}}{{\sin}^{2}}(\theta-\varphi)},$
which is the desired conclusion. ∎
Combining the two propositions above allows us to compute the trace term of (5) for any $X$ with uncorrelated components. For $X$ with non-zero mean, we have the following.
Proposition 2.7.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}\in\mathbb{R}^{2}$ have uncorrelated components, with $\mathbb{E}[X_{1}],\mathbb{E}[X_{2}]$ not necessarily 0. Then, for any $\theta,\varphi\in[0,2\pi)$,
$|\mathbb{E}[R_{\theta}X]-\mathbb{E}[R_{\varphi}X]|^{2}=2\left(\mathbb{E}[X_{1}]^{2}+\mathbb{E}[X_{2}]^{2}\right)(1-\cos(\theta-\varphi)).$
Proof.
$R_{\theta}X-R_{\varphi}X=\begin{bmatrix}X_{1}(\cos\theta-\cos\varphi)-X_{2}(\sin\theta-\sin\varphi)\\ X_{1}(\sin\theta-\sin\varphi)+X_{2}(\cos\theta-\cos\varphi)\end{bmatrix},$
whence by taking expected value and then the square of the Euclidean norm we get
$\left(\mathbb{E}[X_{1}]^{2}+\mathbb{E}[X_{2}]^{2}\right)\left(2-2(\cos\theta\cos\varphi+\sin\theta\sin\varphi)\right)=2\left(\mathbb{E}[X_{1}]^{2}+\mathbb{E}[X_{2}]^{2}\right)(1-\cos(\theta-\varphi))$
as claimed. ∎
Combining Corollary 1.3 with Remark 2.3 and Propositions 2.4, 2.6, and 2.7 with $X_{i}$ replaced by $X_{i}-\mathbb{E}[X_{i}]$, we arrive at the following.
Theorem 2.8.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}\in\mathbb{R}^{2}$ have uncorrelated components. Then for any $\theta,\varphi\in[0,2\pi)$,
$W_{2}(R_{\theta}X,R_{\varphi}X)^{2}\geq 2\left((\mathbb{E}[X_{1}])^{2}+(\mathbb{E}[X_{2}])^{2}\right)(1-\cos(\theta-\varphi))+2(\operatorname{Var}[X_{1}]+\operatorname{Var}[X_{2}])\\ -2\sqrt{(\operatorname{Var}[X_{1}]-\operatorname{Var}[X_{2}])^{2}\cos^{2}(\theta-\varphi)+4\operatorname{Var}[X_{1}]\operatorname{Var}[X_{2}]}$ (7)
2.4. Illustrating the Lower bound for Rotations
Here we give some further perspectives on the lower bound of Theorem 2.8 from various directions. We begin by noticing that the lower bound is equivalent to the Euclidean distance between the angles (for small difference in angles). In particular, for random vectors for which equality holds in 2.8, we see that $W_{2}(R_{\theta}X,R_{\varphi}X)\asymp|\theta-\varphi|$ (that is, there are constants $c_{1},c_{2}\geq 0$ such that $c_{1}|\theta-\varphi|\leq W_{2}(R_{\theta}X,R_{\varphi}X)\leq c_{2}|\theta-\varphi|$) for $|\theta-\varphi|\leq\frac{\pi}{2}$.
Theorem 2.9.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ be a random vector on $\mathbb{R}^{2}$ with uncorrelated components with $m_{i}=\mathbb{E}[X_{i}]$ and $\sigma_{i}^{2}=\operatorname{Var}[X_{i}]$, $i=1,2$. Let $g(X,\theta,\varphi)$ be the lower bound of (7), i.e.,
$g(X,\theta,\varphi)=2(m_{1}^{2}+m_{2}^{2})(1-\cos(\theta-\varphi)+2(\sigma_{1}^{2}+\sigma_{2}^{2})-2\sqrt{(\sigma_{1}^{2}-\sigma_{2}^{2})^{2}\cos^{2}(\theta-\varphi)+4\sigma_{1}^{2}\sigma_{2}^{2}}.$
Then for $|\theta-\varphi|\leq\frac{\pi}{2},$ we have
$\left(\frac{4(m_{1}^{2}+m_{2}^{2})}{\pi^{2}}+\frac{4(\sigma_{1}^{2}-\sigma_{2}^{2})^{2}}{(4+\pi^{2})(\sigma_{1}^{2}+\sigma_{2}^{2})}\right)|\theta-\varphi|^{2}\leq g(X,\theta,\varphi)\leq\left(m_{1}^{2}+m_{2}^{2}+\frac{(\sigma_{1}^{2}-\sigma_{2}^{2})^{2}}{\sigma_{1}\sigma_{2}}\right)|\theta-\varphi|^{2}.$
Proof.
Regarding the upper bound, since $1-\cos x=2\sin^{2}\frac{x}{2}$ and $|\sin z|\leq|z|$ we have
$2(m_{1}^{2}+m_{2}^{2})(1-\cos(\theta-\varphi))\leq(m_{1}^{2}+m_{2}^{2})|\theta-\varphi|^{2}.$
Next, to find the upper bound on the other terms, with $a=\sigma_{1}^{2}=\operatorname{Var}[X_{1}],b=\sigma_{2}^{2}=\operatorname{Var}[X_{2}]$, we define the following function:
$f(x)=2(a+b)-2\sqrt{(a-b)^{2}\cos^{2}x+4ab}$
and notice that $f(0)=0$, whereby for every $0<x<\frac{\pi}{2}$, there exists a $0<\xi<x$ such that $f(x)=f(0)+f^{\prime}(\xi)x$ by the Mean Value Theorem. Now,
$f^{\prime}(\xi)=\frac{(a-b)^{2}2\sin\xi\cos\xi}{\sqrt{(a-b)^{2}\cos^{2}\xi+4ab}}=\frac{(a-b)^{2}\sin(2\xi)}{\sqrt{(a-b)^{2}\cos^{2}\xi+4ab}},$
whence, using that $|\sin(t)|\leq t$ and $|\xi|<|x|$,
$|f(x)|=|f^{\prime}(\xi)x|\leq\frac{(a-b)^{2}|\sin(2\xi)||x|}{2\sqrt{ab}}\leq\frac{(a-b)^{2}|x|^{2}}{\sqrt{ab}}.$
Combining the above estimates with $x=\theta-\varphi$ yields the desired upper bound.
[…]
$2(a+b)-2\sqrt{(a-b)^{2}\cos^{2}x+4ab}=2(a+b)-2\sqrt{(a+b)^{2}-(a-b)^{2}\sin^{2}x}.$
Without loss of generality, we may assume that $a\geq b$, otherwise we can change coordinates of $X$ which will not affect the Wasserstein distance of its rotated copies, hence its lower bound. Then let $A=a+b$ and $B=a-b$, and we have $A>B\geq 0.$ Since $\frac{x}{\sqrt{1+x^{2}}}\leq\sin x$ for $0\leq x\leq\frac{\pi}{2}$, and $\sqrt{1-u^{2}}\leq 1-\frac{1}{2}u^{2}$ for $0\leq u\leq 1,$
$2A-2\sqrt{A^{2}-B^{2}\sin^{2}x}$ $\geq 2A-2\sqrt{A^{2}-B^{2}\frac{x^{2}}{1+x^{2}}}$
$=2A\left(1-\sqrt{1-\left(\frac{B}{A}\frac{x}{\sqrt{1+x}^{2}}\right)^{2}}\right)$
$\geq 2A\left(1-\left(1-\frac{1}{2}\left(\frac{B}{A}\frac{x}{\sqrt{1+x^{2}}}\right)^{2}\right)\right)$
$=\frac{B^{2}}{A}\frac{x}{1+x^{2}}$
$\geq\frac{B^{2}}{A}\frac{x^{2}}{1+\frac{\pi^{2}}{4}}$
$=\frac{B^{2}}{A}\frac{4}{4+\pi^{2}}x^{2}.$
Putting these together, we have
$W_{2}(R_{\theta}X,R_{\varphi}X)^{2}\geq\left(\frac{4(m_{1}^{2}+m_{2}^{2})}{\pi^{2}}+\frac{4(\sigma_{1}^{2}-\sigma_{2}^{2})^{2}}{(4+\pi^{2})(\sigma_{1}^{2}+\sigma_{2}^{2})}\right)|\theta-\varphi|^{2}.$
Corollary 2.10.
When $X$ is a Gaussian, or any random vector that satisfies equality in the lower bound of Theorem 1.2 (hence of Theorem 2.8), one can replace $g(X,\theta,\varphi)$ with $W_{2}(R_{\theta}X,R_{\varphi}X)$ in Theorem 2.9. In particular, for such random vectors, the Wasserstein distance between rotated copies is equivalent to the absolute value of the difference of the rotation angles when $|\theta-\varphi|\leq\frac{\pi}{2}$.
For measures which do not satisfy equality of the lower bound of Theorem 2.8, it is unclear how tight the bound is. We now present some examples in which equality may not be achieved, and illustrate the quality of the lower bound.
Example 2.11.
Let $X$ be uniformly distributed over the unit square $[0,1]\times[0,1]\subset\mathbb{R}^{2}.$ Then $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ where $X_{i}\sim\mathcal{U}[0,1]$ i.i.d., and are thus uncorrelated. We have densities $\mathbbm{1}_{[0,1]}(x_{1})$ and $\mathbbm{1}_{[0,1]}(x_{2})$ with $\mathbb{E}[X_{1}]=\mathbb{E}[X_{2}]=\frac{1}{2},\operatorname{Var}[X_{1}]=\operatorname{Var}[X_{2}]=\frac{1}{{12}}$. Then by Theorem 2.8, for $0\leq|\theta-\varphi|\leq\frac{\pi}{2},$ we have
$W_{2}(R_{\theta}X,R_{\varphi}X)^{2}\geq 2\left(\mathbb{E}[X_{1}]^{2}+\mathbb{E}[X_{2}]^{2}\right)(1-\cos(\theta-\varphi))+2(\operatorname{Var}[X_{1}]+\operatorname{Var}[X_{2}])\\ -4\sqrt{\operatorname{Var}[X_{1}]\operatorname{Var}[X_{2}]}=1-\cos(\theta-\varphi).$
To illustrate the quality of the lower bound in this instance, Figure 1 shows both the actual Wasserstein distance of $X$ and $R_{\theta}X$, the lower bound, and the difference with the lower bound for $\theta\in[0,\frac{\pi}{2}]$. We know that the square does not give equality in the lower bound of 2.8, but nonetheless it appears to be a close approximation in this instance.
Figure 1. Rotation of $[0,1]^{2}$ from Example 2.11. (Left) $W_{2}(X,R_{\theta}X)^{2}$ and the lower bound $1-\cos(\theta)$ for $\theta\in[0,\frac{\pi}{2}]$. (Right) Difference of $W_{2}(X,R_{\theta}X)^{2}-(1-\cos(\theta))$ for $\theta\in[0,\frac{\pi}{2}]$.
Example 2.12.
Let $X$ be uniformly distributed over the centered unit square $[-\frac{1}{2},\frac{1}{2}]^{2}\subset\mathbb{R}^{2}$. Then $\mathbb{E}[X_{i}]=0$ and $\operatorname{Var}[X_{i}]=\frac{1}{12}$ for $i=1,2$. Then the lower bound of Theorem 2.8 is 00, and we only obtain the trivial bound $W_{2}(R_{\theta}X,R_{\varphi}X)^{2}\geq 0$. Figure 2 shows the Wasserstein distance of $X$ and $R_{\theta}X$ for $\theta\in[0,\frac{\pi}{2}]$. Note that due to the fact that the centered square has rotational symmetry at multiples of $\frac{\pi}{2}$, we see that the Wasserstein distance is increasing on $[0,\frac{\pi}{4}]$ and decreasing on $[\frac{\pi}{4},\frac{\pi}{2}]$ as expected.
Figure 2. Rotation of $[-\frac{1}{2},\frac{1}{2}]^{2}$ from Example 2.12. Plotted is $W_{2}(X,R_{\theta}X)^{2}$ for $\theta\in[0,\frac{\pi}{2}]$.
Example 2.13.
Let $X$ be the uniformly distributed on the rectangle $[0,2]\times[0,1]\subset\mathbb{R}^{2}$ with densities $\frac{1}{2}\mathbbm{1}_{[0,2]}({x_{1}}),\mathbbm{1}_{[0,1]}(x_{2})$ of its independent components $X_{1},X_{2}$. Then $\mathbb{E}[X_{1}]=1,\mathbb{E}[X_{2}]=\frac{1}{2},\operatorname{Var}[X_{1}]=\frac{1}{3},\operatorname{Var}[X_{2}]=\frac{1}{12}$ and the lower bound of Theorem 2.8 is
$W_{2}(R_{\theta}X,R_{\varphi}X)^{2}\geq\frac{5}{2}(1-\cos(\theta-\varphi))+\left(\frac{5}{6}-\sqrt{\frac{\cos^{2}(\theta-\varphi)}{4}+\frac{4}{9}}\right).$ (8)
To illustrate the quality of the lower bound in this instance, Figure 3 shows both the actual Wasserstein distance of $X$ and $R_{\theta}X$, the lower bound, and the difference with the lower bound for $\theta\in[0,\frac{\pi}{2}]$. We know that the rectangle does not give equality in the lower bound of 2.8, but nonetheless it appears to be a relatively close approximation in this instance.
Figure 3. Rotation of $[0,2]\times[0,1]$ from Example 2.13. (Left) $W_{2}(X,R_{\theta}X)^{2}$ and the lower bound of (8) for $\theta\in[0,\frac{\pi}{2}]$. (Right) Difference of $W_{2}(X,R_{\theta}X)^{2}$ and the lower bound of (8) for $\theta\in[0,\frac{\pi}{2}]$.
We end this section with a proposition that shows that for random vectors with independent, zero-mean components, the lower bound of Theorem 2.8 can be achieved by an explicit map, which turns out to be that mentioned in Dowson and Landau’s result (Theorem 1.2).
Proposition 2.14.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ where $X_{1}$ and $X_{2}$ are independent random variables with mean 0 and $\operatorname{Var}[X_{1}]=\mathbb{E}[X_{1}^{2}]=a$, $\operatorname{Var}[X_{2}]=\mathbb{E}[X_{2}^{2}]=b$, and let $Y=R_{\theta}X$. Let $T:=\Sigma_{X}^{-1}(\Sigma_{X}\Sigma_{R_{\theta}X})^{\frac{1}{2}}$. Then
$T=\frac{1}{\sqrt{4ab+(a-b)^{2}\cos^{2}\theta}}\begin{bmatrix}b+a\cos^{2}\theta+b\sin^{2}\theta&(a-b)\cos\theta\sin\theta\\ (a-b)\cos\theta\sin\theta&a+b\cos^{2}\theta+a\sin^{2}\theta\end{bmatrix}$
and $\det(T)=1.$ Moreover,
$\mathbb{E}|TX-X|^{2}=2(a+b)-2\sqrt{4ab+(a-b)^{2}\cos^{2}\theta}\leq W_{2}(X,R_{\theta}X)^{2}.$
In particular, the map $T$ attains the lower bound for the Wasserstein distance between $X$ and $R_{\theta}X$ in (3).
Proof.
Clearly $\Sigma_{X}=\textnormal{diag}(a,b)$ is invertible. Thus, $T=\Sigma_{X}^{-1}(\Sigma_{X}\Sigma_{R_{\theta}X})^{\frac{1}{2}}$ is such that $\mathbb{E}|TX-X|^{2}=\textnormal{tr}[\Sigma_{X}+\Sigma_{R_{\theta}X}-2(\Sigma_{X}\Sigma_{R_{\theta}X})^{\frac{1}{2}}]$ by Theorem 1.2 (although one can verify this directly as well). It remains to compute $T$.
Noting that
$\Sigma_{X}=\textnormal{diag}(a,b),\qquad\Sigma_{R_{\theta}X}=\begin{bmatrix}a\cos^{2}\theta+b\sin^{2}\theta&(a-b)\cos\theta\sin\theta\\ (a-b)\cos\theta\sin\theta&b\cos^{2}\theta+a\sin^{2}\theta\end{bmatrix},$
we compute $\textnormal{tr}(\Sigma_{X}\Sigma_{R_{\theta}X})=(a^{2}+b^{2})\cos^{2}\theta+2ab\sin^{2}\theta$ and $\det(\Sigma_{X}\Sigma_{R_{\theta}X})=a^{2}b^{2}$, and thus by Lemma 2.5,
$\textnormal{tr}(\Sigma_{X}\Sigma_{R_{\theta}X})+2\sqrt{\det(\Sigma_{X}\Sigma_{R_{\theta}X})}=(a^{2}+b^{2})\cos^{2}\theta+2ab\sin^{2}\theta+2ab=4ab+(a-b)^{2}\cos^{2}\theta.$
Therefore,
$(\Sigma_{X}\Sigma_{R_{\theta}X})^{\frac{1}{2}}=\frac{1}{\sqrt{4ab+(a-b)^{2}\cos^{2}\theta}}\begin{bmatrix}a(a\cos^{2}\theta+b\sin^{2}\theta+b)&a(a-b)\cos\theta\sin\theta\\ b(a-b)\cos\theta\sin\theta&b(b\cos^{2}\theta+a\sin^{2}\theta+a)\end{bmatrix}.$
Multiplying this matrix by $\Sigma_{X}^{-1}=\textnormal{diag}(\frac{1}{a},\frac{1}{b})$ gives the desired formula for $T$. ∎
The utility of the above proposition is that, even though the Wasserstein distance between $X$ and $R_{\theta}X$ is difficult to ascertain in general, we can easily form an explicit map $T$ that achieves the lower bound for the Wasserstein distance.
2.5. Compositions of affine maps
In this section we derive an upper bound for the Wasserstein distance between probability measures corresponding to a random vector and a composition of affine transformations. Recall that $T_{\alpha}$, $S_{\lambda}$, and $R_{\theta}$ represent translation, scaling, and rotation maps, respectively. We consider compositions of these, which can be utilized to understand a larger parametrized subset of diffeomorphisms.
Theorem 2.15.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ be a random vector in $\mathbb{R}^{2}$ with uncorrelated components. Let $\alpha\in\mathbb{R}$, $\lambda\in\mathbb{R}^{2}$ with $\lambda_{1},\lambda_{2}\neq 0$, and $\theta\in[0,2\pi)$. Suppose also that equality holds in Theorem 2.8 for $W_{2}(R_{\theta}S_{\lambda}X,S_{\lambda}X)$. Then,
$W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)\leq|\alpha|+\sqrt{(\lambda_{1}-1)^{2}\mathbb{E}[X_{1}^{2}]+(\lambda_{2}-1)^{2}\mathbb{E}[X_{2}^{2}]}\\ +\left\{2(\lambda_{1}^{2}\mathbb{E}[X_{1}]^{2}+\lambda_{2}^{2}\mathbb{E}[X_{2}]^{2})(1-\cos\theta)+2(\lambda_{1}^{2}\operatorname{Var}[X_{1}]+\lambda_{2}^{2}\operatorname{Var}[X_{2}])\right.\\ \left.-2\sqrt{(\lambda_{1}^{2}\operatorname{Var}[X_{1}]-\lambda_{2}^{2}\operatorname{Var}[X_{2}])^{2}\cos^{2}\theta+4\lambda_{1}^{2}\lambda_{2}^{2}\operatorname{Var}[X_{1}]\operatorname{Var}[X_{2}]}\right\}^{\frac{1}{2}}$
Proof.
Note that
$W_{2}(T_{\alpha}R_{\theta}S_{\lambda}X,X)$ $\leq W_{2}(S_{\lambda}X,X)+W_{2}(R_{\theta}S_{\lambda}X,S_{\lambda}X)+W_{2}(T_{\alpha}R_{\theta}S_{\lambda}X,R_{\theta}S_{\lambda}X)$
$=:I_{1}+I_{2}+I_{3}.$
By Proposition 2.1, $I_{3}=|\alpha|$. By Proposition 2.2 $I_{1}=\sqrt{(\lambda_{1}-1)^{2}\mathbb{E}[X_{1}^{2}]+(\lambda_{2}-1)^{2}\mathbb{E}[X_{2}^{2}]}.$
Finally, to estimate $I_{2}$, we apply Theorem 2.8 with $\varphi=0$ and $X$ replaced with $S_{\lambda}X$. Noting that $S_{\lambda}X=\begin{bmatrix}\lambda_{1}X_{1}\\ \lambda_{2}X_{2}\end{bmatrix}$ and $\mathbb{E}[(S_{\lambda}X)_{i}]=\lambda_{i}\mathbb{E}[X_{i}]$, $i=1,2$, we have
$I_{2}^{2}=2\left(\lambda_{1}^{2}(\mathbb{E}[X_{1}])^{2}+\lambda_{2}^{2}(\mathbb{E}[X_{2}])^{2}\right)(1-\cos\theta)+2(\lambda_{1}^{2}\operatorname{Var}[X_{1}]+\lambda_{2}^{2}\operatorname{Var}[X_{2}])\\ -2\sqrt{(\lambda_{1}^{2}\operatorname{Var}[X_{1}]-\lambda_{2}^{2}\operatorname{Var}[X_{2}])^{2}\cos^{2}(\theta)+4\lambda_{1}^{2}\lambda_{2}^{2}\operatorname{Var}[X_{1}]\operatorname{Var}[X_{2}]}.$
Combining the estimates above yields the result. ∎
Theorem 2.16.
With the setup of Theorem 2.15, suppose that equality does not necessarily hold in Theorem 2.8. Then,
$W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)\leq|\alpha|+\sqrt{(\lambda_{1}-1)^{2}\mathbb{E}[X_{1}^{2}]+(\lambda_{2}-1)^{2}\mathbb{E}[X_{2}^{2}]}\\ \quad+\left\{2(\lambda_{1}^{2}\mathbb{E}[X_{1}]^{2}+\lambda_{2}^{2}\mathbb{E}[X_{2}]^{2})(1-\cos\theta)+2(\lambda_{1}^{2}\operatorname{Var}[X_{1}]+\lambda_{2}^{2}\operatorname{Var}[X_{2}])\right.\\ \left.+2\sqrt{(\lambda_{1}^{2}\operatorname{Var}[X_{1}]-\lambda_{2}^{2}\operatorname{Var}[X_{2}])^{2}\cos^{2}\theta+4\lambda_{1}^{2}\lambda_{2}^{2}\operatorname{Var}[X_{1}]\operatorname{Var}[X_{2}]}.\right\}^{\frac{1}{2}}$
The difference in the bounds of Theorems 2.15 and 2.16 is merely the $+$ sign in the final term by using the upper bound of Dowson and Landau [5] that $\mathbb{E}|X-Y|^{2}\leq\textnormal{tr}[\Sigma_{X}+\Sigma_{Y}+2(\Sigma_{X}\Sigma_{Y})^{\frac{1}{2}}]$. This upper bound is typically far from optimal as can be seen in the examples above given that the lower bound is such a close approximation. Nonetheless it does give a qualitative upper bound for composition maps which is small for sufficiently small angles and dilations.
Remark 2.17.
Unfortunately, changing the order of the maps in the composition does not allow the proof of the theorem above to go through. In particular, application of Theorem 2.8 requires that the random vector in question has uncorrelated components, which would not be true if the rotation was applied first followed by scaling.
Example 2.18.
Consider $X$ to be drawn from the uniform measure on the unit disk centered at the origin in $\mathbb{R}^{2}$, so here $(\Omega,F,P)=([0,2\pi)\times[0,1],\mathcal{B}([0,2\pi)\times[0,1]),\frac{d\theta}{2\pi}\times dr)$, $X=\begin{bmatrix}r\cos\theta\\ r\sin\theta\end{bmatrix}$,
$\mathbb{E}[X_{1}]=\frac{1}{2\pi}\int_{0}^{2\pi}\int_{0}^{1}r\cos\theta drd\theta=0,\qquad\mathbb{E}[x_{1}]=\frac{1}{2\pi}\int_{0}^{2\pi}\int_{0}^{1}r\sin\theta drd\theta=0,$
$\operatorname{Var}[X_{1}]=\frac{1}{2\pi}\int_{0}^{2\pi}\int_{0}^{1}r^{2}\cos^{2}\theta drd\theta=\frac{1}{6},\qquad\operatorname{Var}[X_{2}]=\frac{1}{2\pi}\int_{0}^{2\pi}\int_{0}^{1}r^{2}\sin^{2}\theta drd\theta=\frac{1}{6},$
and
[…]
hence $X$ has uncorrelated components.
This distribution satisfies the elliptically contoured condition definition, and thus achieves the lower bound of Corollary 1.3, and therefore we may apply Theorem 2.15 to see that
$W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)\leq|\alpha|+\frac{1}{\sqrt{6}}|\lambda-\mathbf{1}|+\left(\frac{1}{3}|\lambda|^{2}-\frac{1}{3}\sqrt{(\lambda_{1}^{2}-\lambda_{2}^{2})^{2}\cos^{2}\theta+4\lambda_{1}^{2}\lambda_{2}^{2}}\right)^{\frac{1}{2}},$ (9)
[…]
Figure 4 illustrates this bound for 50 samples on the unit disk, 10 angles $\theta\in[0,\frac{\pi}{2}]$, $\alpha=\begin{bmatrix}\frac{1}{2}\\ \frac{1}{2}\end{bmatrix}$, and $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$.
Figure 4. $X$ drawn from the uniform distribution on the unit disk in $\mathbb{R}^{2}$ as in Example 2.18. (Left) $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$ and the upper bound of (9) for $\theta\in[0,2\pi]$ with $\alpha=\begin{bmatrix}\frac{1}{2}\\ \frac{1}{2}\end{bmatrix}$, $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$. (Right) Difference of the upper bound and $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$.
3. Applications to 1-dimensional submanifolds
Now we will present a number of illustrative examples of Wasserstein distance under composition of scaling/rotation/translation for uniform distributions on a unit circle, line segment, and other curves. Such examples naturally arise in the context of alphabet letters recognition subject to affine distortions, especially when considering small perturbations from the actual objects.
Example 3.1.
Consider a uniform measure on a unit circle, which can be realized as follows $(\Omega,F,P)=([0,2\pi],\mathcal{B}([0,2\pi]),\frac{dt}{2\pi})$, where $\frac{dt}{2\pi}$ is the Lebesgue measure normalized to be a probability measure on $[0,2\pi]$, i.e., the uniform distribution with density $\frac{1}{2\pi}\mathbbm{1}_{[0,2\pi]}(t).$ Consider a random vector
$X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}:\Omega\to S^{1},\quad\textnormal{via}\quad X(t)=\begin{bmatrix}X_{1}(t)\\ X_{2}(t)\end{bmatrix}=\begin{bmatrix}\cos t\\ \sin t\end{bmatrix},t\in[0,2\pi].$
Then $X(t)$ is uniformly distributed over the unit circle $S^{1}$.
Note that since $\int_{0}^{2\pi}\cos tdt=\int_{0}^{2\pi}\sin tdt=0$, we have $\mathbb{E}[X]=0.$ Similarly, $\mathbb{E}[X_{1}X_{2}]=0$, as $\int_{0}^{2\pi}\cos t\sin tdt=0$, hence $X_{1}$ and $X_{2}$ are uncorrelated. Finally,
$\operatorname{Var}[X_{1}]=\mathbb{E}[X_{1}^{2}]=\frac{1}{2\pi}\int_{0}^{2\pi}\cos^{2}tdt=\frac{1}{2}=\frac{1}{2\pi}\int_{0}^{2\pi}\sin^{2}tdt=\mathbb{E}[X_{2}^{2}]=\operatorname{Var}[X_{2}].$
Note that in this instance, $X$ satisfies the criteria for equality of the lower bound of Corollary 1.3 (see Gelbrich [7]). Consequently, the conclusion of Theorem 2.15 yields
$W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)\leq|\alpha|+\frac{1}{\sqrt{2}}|\lambda-\mathbf{1}|+\left(|\lambda|^{2}-\sqrt{(\lambda_{1}^{2}-\lambda_{2}^{2})^{2}\cos^{2}\theta+4\lambda_{1}^{2}\lambda_{2}^{2}}\right)^{\frac{1}{2}}.$ (10)
Note that this bound is more useful for small values of $\alpha,\theta,\lambda$, but nonetheless gives an explicit upper bound in terms of these parameters. Figure 5 illustrates a simple case of this bound for 50 samples from the uniform distribution on $S^{1}$ and 50 values of $\theta\in[0,2\pi]$ for $\alpha=0$ and $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$. In this case, we see equality at the endpoints as expected because there is no contribution from $|\alpha|$. Figure 6 uses the same setup but with $\alpha=\begin{bmatrix}-\frac{1}{2}\\ 1\end{bmatrix}$. Here, we see that the translation yields a worse upper bound as expected, although the characteristic of the bound is qualitatively similar.
Figure 5. $X$ drawn from the uniform distribution on the unit circle in $\mathbb{R}^{2}$ as in Example 3.1. (Left) $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$ and the upper bound of (10) for $\theta\in[0,2\pi]$ with $\alpha=0$, $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$. (Right) Difference of the upper bound and $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$.
Figure 6. $X$ drawn from the uniform distribution on the unit circle in $\mathbb{R}^{2}$ as in Example 3.1. (Left) $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$ and the upper bound of (10) for $\theta\in[0,2\pi]$ with $\alpha=\begin{bmatrix}-\frac{1}{2}\\ 1\end{bmatrix}$ and $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$. (Right) Difference of the upper bound and $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$.
Example 3.2.
Let $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}=\begin{bmatrix}X_{1}\\ 0\end{bmatrix}$, where $X$ is uniformly distributed on $[-\frac{1}{2},\frac{1}{2}]$ with density $\mathbbm{1}_{[-\frac{1}{2},\frac{1}{2}]}(t).$ Then $\mathbb{E}[X_{1}]=\mathbb{E}[X_{2}]=0$, $\operatorname{Var}[X_{1}]=\frac{1}{12}$, $\operatorname{Var}[X_{2}]=0$. One can readily verify in this case that $\mathbb{E}[|X-R_{\theta}X|^{2}=\mathbb{E}[X_{1}^{2}(2(1-\cos\theta))]=\frac{1}{6}(1-\cos\theta)$. Additionally, one can compute the trace quantity in the lower bound of (3) directly and see that it is $\frac{1}{6}(1-\cos\theta)$. Therefore, we may use Theorem 2.15, which implies that
$W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)\leq|\alpha|+\frac{1}{\sqrt{12}}|\lambda_{1}-1|+\frac{1}{\sqrt{3}}|\lambda_{1}|^{2}\sin\frac{\theta}{2}.$ (11)
Note that in deriving the bound above from Theorem 2.15. we have used the fact that $\sqrt{1-\cos\theta}=\sqrt{2}\sin\frac{\theta}{2}$.
Since $X$ is drawn from a 1-dimensional subset of $\mathbb{R}^{2}$, it makes sense that only dilation in the $x$-direction will impact the upper bound (11) (since dilation is the first operation). The upper bound here is evidently good for small translations and dilations as in Example 3.1, but will not be for large translation or dilation. Figure 7 shows a simple example of this upper bound for $\alpha=\begin{bmatrix}-.25\\ .1\end{bmatrix}$ and $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$, in which 30 samples are drawn from $\mathbbm{1}_{[-\frac{1}{2},\frac{1}{2}]}$ and 20 values of $\theta\in[0,\frac{\pi}{2}]$ are used.
Figure 7. $X$ drawn from the uniform distribution on $[-\frac{1}{2},\frac{1}{2}]$ in $\mathbb{R}^{2}$ as in Example 3.2. (Left) $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$ and the upper bound of (11) for $\theta\in[0,2\pi]$ with $\alpha=\begin{bmatrix}-.25\\ .1\end{bmatrix}$ and $\lambda=\begin{bmatrix}2\\ \frac{1}{2}\end{bmatrix}$. (Right) Difference of the upper bound and $W_{2}(T_{\alpha}\circ R_{\theta}\circ S_{\lambda}X,X)$.
3.1. Toward handwriting analysis
In the following examples we illustrate construction of uniform distributions over shapes represented by random curves with uncorrelated components, whereby satisfying the assumptions for Wasserstein distance formulas developed in this paper.
We focus on deriving random vectors $X\in\mathbb{R}^{2}$ which could represent handwritten letters with the notion that these could be used in conjunction with the bounds in the previous sections to better understand Wasserstein distances between images in handwriting datasets like MNIST [10]. This has potential utility in a variety of avenues including clustering (e.g., [3]) or manifold learning. In particular, several recent works have explored the use of optimal transport in manifold learning (i.e., learning submanifolds of $W_{2}(\mathbb{R}^{d})$) [4, 9, 12, 13, 14, 20].
Utilizing the constructions below as well as template deformations such as the composition maps in Section 2.5, one could develop a synthetic handwritten image dataset for testing in machine learning applications. A preliminary example of this was done for handwritten 00’s in [9].
Example 3.3 (Square).
Let $(\Omega,F,P)=([0,4],\mathcal{B}([0,4]),\frac{dt}{4})$, and consider
$X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix},\quad\textnormal{where}\quad X(t)=\begin{bmatrix}X_{1}(t)\\ X_{2}(t)\end{bmatrix},\quad 0\leq t\leq 4$
with
$X_{1}(t)=\begin{cases}\frac{1}{2}-t,&0\leq t<1\\ -\frac{1}{2},&1\leq t<2\\ t-\frac{5}{2},&2\leq t<3\\ \frac{1}{2},&3\leq t\leq 4,\end{cases}\qquad X_{2}(t)=\begin{cases}-\frac{1}{2},&0\leq t<1\\ t-\frac{3}{2},&1\leq t<2\\ \frac{1}{2},&2\leq t<3\\ \frac{7}{2}-t,&3\leq t\leq 4.\end{cases}$
It can be readily verified that $\mathbb{E}[X_{1}]=\mathbb{E}[X_{2}]=\mathbb{E}[X_{1}X_{2}]=0$.
Figure 8. (Left) $X_{1}$ and $X_{2}$ vs. $t$. (Right) $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ forming a square.
Example 3.4 (Letter T).
Let $(\Omega,F,P)$ and $X$ be as in Example 3.3 with
$X_{1}(t)=\begin{cases}1-t,&0\leq t\leq 2\\ 0,&2<t\leq 4,\end{cases},\qquad X_{2}(t)=\begin{cases}\frac{1}{2},&0\leq t\leq 2\\ t-\frac{7}{2},&2<t\leq 4.\end{cases}$
Again, it is readily verified that $\mathbb{E}[X_{1}]=\mathbb{E}[X_{2}]=\mathbb{E}[X_{1}X_{2}]=0.$
Figure 9. (Left) $X_{1}$ and $X_{2}$ vs. $t$. (Right) $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ forming the letter T.
Example 3.5 (Letter A).
Let $(\Omega,F,P)=([0,6],\mathcal{B}([0,6]),\frac{dt}{6})$, and consider $X$ defined as $X(t)=\begin{bmatrix}X_{1}(t)\\ X_{2}(t)\end{bmatrix}$, $t\in[0,6]$, where
$X_{1}(t)=\begin{cases}2-t,&0\leq t\leq 4\\ t-5,&4<t\leq 6,\end{cases}\qquad X_{2}(t)=\begin{cases}t-1,&0\leq t\leq 2\\ 3-t,&2<t\leq 4\\ 0,&4<t\leq 6.\end{cases}$
Figure 10. (Left) $X_{1}$ and $X_{2}$ vs. $t$. (Right) $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ forming the letter A.
Example 3.6 (Letter C).
Let $(\Omega,F,P)=([\frac{\pi}{2},\frac{3\pi}{2},\mathcal{B}([\frac{\pi}{2},\frac{3\pi}{2}]),\frac{dt}{\pi})$, and let
$X(t)=\begin{bmatrix}X_{1}(t)\\ X_{2}(t)\end{bmatrix}=\begin{bmatrix}\cos t+\frac{2}{\pi}\\ \sin t\end{bmatrix},\quad t\in\left[\frac{\pi}{2},\frac{3\pi}{2}\right].$
$X$ is the semicircle of radius 1 centered at $(\frac{2}{\pi},0)$, and one can readily check that $\mathbb{E}[X_{1}]=\mathbb{E}[X_{2}]=\mathbb{E}[X_{1}X_{2}]=0$.
Figure 11. (Left) $X_{1}$ and $X_{2}$ vs. $t$. (Right) $X=\begin{bmatrix}X_{1}\\ X_{2}\end{bmatrix}$ forming the letter C.
We have shown in this subsection how several letters of the alphabet can be formed by random vectors in $\mathbb{R}^{2}$ that satisfy the criteria in previous sections which allow us to determine Wasserstein distance between these letters and certain deformations of them. In particular, Theorem 2.15 applied to these examples allows one to understand the Wasserstein distance between an idealized letter A (for instance) and a dilated, rotated, and translated copy of it.
Acknowledgements
KH was partially supported by a Research Enhancement Program grant from the College of Science at the University of Texas at Arlington. Research was sponsored by the Army Research Office and was accomplished under Grant Number W911NF-23-1-0213. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.